# 0000-0003-0424-585X

### v4

@@ 1 Introduction
Standard economic models of optimal growth typically assume an infinite planning horizon. This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey1928). While this framework functions well in growing economies, it raises deep ethical concerns when applied to stagnant or declining economies with resource constraints. In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (DASGUPTA2019). By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (PARFIT).
@@
This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005). Unlike standard utilitarianism, which sums utilities regardless of their sign, CLU introduces a normative threshold of well-being (the critical level). An individual’s existence contributes to social welfare only if their utility exceeds this critical level; otherwise, it is considered a negative contribution. By applying this ethical principle to a dynamic setting, we pose a question that is often excluded from standard growth theory: Under what economic conditions is it ethically optimal for a dynasty to continue, and when is its planned termination the morally justified conclusion of history?
@@
Methodologically, our approach establishes a structural isomorphism between static population ethics and dynamic growth theory. In the tradition of optimal population theory initiated by Samuelson1975, a social planner typically chooses the optimal population size $N$ to maximize social welfare under resource constraints. In our dynamic framework, we reinterpret this problem by viewing the number of generations $N$ (longevity) as the choice variable. This mapping reveals that the “Repugnant Conclusion” (a large population with low quality of life) and “Infinite Misery” (an infinite dynasty with low quality of life) are two sides of the same ethical coin. Consequently, our finite-horizon solution implies a conscious decision to cease reproduction, rather than a failure of time consistency. It represents an altruistic decision to prevent the creation of future generations whose lives would fall below the critical level of well-being.
@@
We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (AsheimZuber2014). RDU prevents the conclusion by strictly discounting the value of additional lives. However, this approach can lead to the opposite extreme—an “anti-population bias” or “reverse repugnant conclusion”—where a small population with high utility is preferred over a larger population that is also well-off. Furthermore, in a dynamic context, RDU implies that the welfare weight of future generations depends on their relative rank in history. This dependency destroys the recursive structure of the optimization problem and leads to time inconsistency. In contrast, CLU provides a transparent, absolute threshold for “lives worth living” that preserves both intergenerational equity (for those above the threshold) and the tractability of dynamic programming.
@@ 2.1 Critical-Level Preferences
Before describing the production technology, we first specify the social planner’s objective function. We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (BlackorbyDonaldson1984; BBD2005). Let $\Upsilon(c)$ denote the lifetime utility (well-being) of an individual consuming a constant level of resources $c$. The planner aims to determine the consumption path and the planning horizon (number of generations) to maximize the following social welfare function:
@@
For analytical tractability and to capture the essential trade-offs, we employ a logarithmic utility function of the form $\Upsilon(c)=\log c+\alpha$. This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus1992; Stern2007), which facilitates comparisons with standard benchmarks. Under this specification, the contribution of generation $t$ to the population value becomes:
@@ 3.1 AK setting
The AK model, which was formally developed by Frankel, is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $N\to\infty$ is an optimal policy. The optimal consumption path (10) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
@@
For the sake of the analysis, let us take the derivative with respect to N{N}.
@@ 4 Concluding Remarks
Methodologically, our approach ensures time consistency through the recursive structure of dynamic programming. Unlike ad hoc termination rules, the finite horizon derived here is the result of a consistent optimization process where the continuation value falls below the critical threshold. We acknowledge, however, that our model relies on a logarithmic utility function and a fixed population size per generation to maintain analytical tractability. Introducing more general preferences (such as CRRA utility), endogenous population size, or environmental constraints would likely preclude closed-form solutions and require numerical approaches. Exploring these complex dynamics remains a promising avenue for future research.

### v5

@@ 1 Introduction
Standard economic models of optimal growth typically assume an infinite planning horizon. This assumption, justified by the mathematical convenience of asymptotic stability or the ethical stance of treating all future generations impartially, implicitly imposes a “perpetual existence” constraint on the dynasty (Ramsey, 1928). While this framework functions well in growing economies, it raises deep ethical concerns when applied to stagnant or declining economies with resource constraints. In such settings, the standard discounted utilitarian approach can lead to a “dismal” long-run equilibrium where consumption approaches zero asymptotically (Dasgupta, 2019). By enforcing indefinite survival, these models may inadvertently justify the perpetuation of lives that are barely worth living, or worse, lives filled with misery—a dynamic analogue to Parfit’s “Repugnant Conclusion” in population ethics (Parfit, 1986).
@@
This paper revisits the problem of dynastic planning through the lens of Critical-Level Utilitarianism (CLU) (Blackorby and Donaldson, 1984; Blackorby et al., 2005). Unlike standard utilitarianism, which sums utilities regardless of their sign, CLU introduces a normative threshold of well-being (the critical level). An individual’s existence contributes to social welfare only if their utility exceeds this critical level; otherwise, it is considered a negative contribution. By applying this ethical principle to a dynamic setting, we pose a question that is often excluded from standard growth theory: Under what economic conditions is it ethically optimal for a dynasty to continue, and when is its planned termination the morally justified conclusion of history?
@@
Methodologically, our approach establishes a structural isomorphism between the spatial and temporal dimensions of optimal population theory. In the tradition initiated by Samuelson (1975), dynamic models of optimal population typically assume an infinite planning horizon and focus on choosing the optimal population growth rate (i.e., the spatial size of cohorts) to maximize social welfare. Recently, Renström and Spataro (2011, 2019) extended this framework to (Relative) CLU. Yet, across this literature, the assumption of perpetual existence remains unquestioned. Our paper complements this literature by inverting the paradigm: rather than optimizing the horizontal size of each generation over an infinite horizon, we fix the cohort size and treat the vertical length of the dynasty—the number of generations $N$ (longevity)—as the endogenous choice variable. This mapping reveals that the “Repugnant Conclusion” (a large population with low quality of life) and “Infinite Misery” (an infinite dynasty with low quality of life) are two sides of the same ethical coin. Consequently, our finite-horizon solution implies a conscious decision to cease reproduction, rather than a failure of time consistency. It represents an altruistic decision to prevent the creation of future generations whose lives would fall below the critical level of well-being.
@@
We acknowledge that other axiomatic frameworks, such as Rank-Discounted Utilitarianism (RDU), have been proposed to avoid the Repugnant Conclusion (Asheim and Zuber, 2014). RDU prevents the conclusion by strictly discounting the value of additional lives. However, this approach can lead to the opposite extreme—an “anti-population bias” or “reverse repugnant conclusion”—where a small population with high utility is preferred over a larger population that is also well-off. Furthermore, in a dynamic context, RDU implies that the welfare weight of future generations depends on their relative rank in history. This dependency destroys the recursive structure of the optimization problem and leads to time inconsistency. In contrast, CLU provides a transparent, absolute threshold for “lives worth living” that preserves both intergenerational equity (for those above the threshold) and the tractability of dynamic programming.
@@ 2.1 Critical-Level Preferences
Before describing the production technology, we first specify the social planner’s objective function. We posit a social planner whose objective is to maximize the population value based on Critical-Level Utilitarianism (CLU) (Blackorby and Donaldson, 1984; Blackorby et al., 2005). Let $\Upsilon(c)$ denote the lifetime utility (well-being) of an individual consuming a constant level of resources $c$. The planner aims to determine the consumption path and the planning horizon (number of generations) to maximize the following social welfare function:
@@
For analytical tractability and to capture the essential trade-offs, we employ a logarithmic utility function of the form $\Upsilon(c)=\log c+\alpha$. This functional form is widely employed in optimal growth literature, including climate change economics (Nordhaus, 1992; Stern, 2007), which facilitates comparisons with standard benchmarks. Under this specification, the contribution of generation $t$ to the population value becomes:
@@ 3.1 AK setting
The AK model, which was formally developed by Frankel (1962), is one of the simplest fundamental models of endogenous growth. Here, we employ this production model to study the population value function that discounts future generations to determine whether $N\to\infty$ is an optimal policy. The optimal consumption path (10) for the AK setting with future discounting ($\theta=1$, $\beta<1$) may be specified as follows:
@@
For the sake of the analysis, let us treat the planning horizon $N$ as a continuous variable and take the derivative with respect to $N$.
@@ 4 Concluding Remarks
Methodologically, our approach ensures time consistency through the recursive structure of dynamic programming. [^7] Formally, suppose that at some intermediate period $t<N^{*}$, the planner re-optimizes and chooses a different horizon $N^{\#}\neq N^{*}$ (and a corresponding consumption path) that yields a higher continuation value than the original plan. This would imply that the original total value could have been improved by switching to this new path at $t$, contradicting the optimality of the original plan $(c_{0}^{*},\dots,c_{N^{*}}^{*})$. Therefore, the optimal horizon $N^{*}$ determined at $t=0$ remains optimal for any subsequent generation $t$. Unlike ad hoc termination rules, the finite horizon derived here is the result of a consistent optimization process where the continuation value falls below the critical threshold. We acknowledge, however, that our model relies on a logarithmic utility function and a fixed population size per generation to maintain analytical tractability. Introducing more general preferences (such as CRRA utility), endogenous population size, or environmental constraints would likely preclude closed-form solutions and require numerical approaches. Exploring these complex dynamics remains a promising avenue for future research.